OpenStudy (anonymous):

You are driving along the highway at 27.8 m/s. An emergency vehicle, driving directly toward you (head-on) at 33.3 m/s in the opposite lane, is blaring its siren. The siren's main tone has a frequency of 700 Hz when the emergency vehicle is at rest. What frequency do you hear from the siren? The speed of sound is 343 m/s.

OpenStudy (shamim):

Required formula is f'=(v+v0)f/(v-vs)

OpenStudy (shamim):

Given v=343m/s v0=27.8m/s vs=33.3m/s f=700hz f'=?